OpenStudy (anonymous):

Identify the 33rd term of the arithmetic sequence 9, 7 1/2, 6 Question 3 options: 1) -40 1/2 2) -39 3) -7 1/2 4) -7

OpenStudy (dumbcow):

-39

OpenStudy (angela210793):

d=-3/2 y33=y1+(n-1)*d=9+32*(-3/2)=9+-48=-39

OpenStudy (anonymous):

-39